OpenStudy (anonymous):

Solve using elimination method. -3x+2y=3 4x+3y=-4

OpenStudy (anonymous):

multiply the first equation by 3/2 and subtract them you will get just one variable

OpenStudy (anonymous):

Whats the solution?

OpenStudy (anonymous):

nvm its (-1,0)

OpenStudy (anonymous):

OK they mention me not to right answers

OpenStudy (anonymous):

\[-3x+2y=3\\4x+3y=-4\]multiple the first raw by 4 and the second one by 3, then we got:\[-12x+8y=12\\12x+9y=-12\]add se first raw to the second one\[-12x+8y=12\\16y=0\]replacing this value at the first raw\[-12x+8\cdot0=12\\y=0\]\[\boxed{\boxed{x=-1~~and~~y=0}}\]\[(-1,0)\]